# Neural Configuration-Space Barriers for Manipulation Planning and Control

### v1

@@ VII-A Governor-based Trajectory Tracking
VII-B Distributionally Robust Neural CDF Barrier
@@
where $\alpha_{h}$ is a class $\mathcal{K}_{\infty}$ function.
@@
$\mathbf{u}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m}}\|\mathbf{u}-\bar{\mathbf{u}}(\mathbf{q},t)\|^{2}$ (17)
s.t. $\;\text{CBC}(\mathbf{q},\mathbf{u},t)\geq 0.$
However, the CDF barrier $h(\mathbf{q},t)$ and its gradients are not known exactly but rather are approximated through the neural CDF barrier $\hat{h}(\mathbf{q},t)$ in (8). Due to uncertainties introduced by the neural network model and the point-cloud data, $\hat{h}$ cannot be treated as an exact representation of $h$. Instead, we work with samples of the associated uncertainty vector $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$, defined in (16), which capture the variations in $\hat{h}$, its gradient, and its time derivative. The practical approach of obtaining such samples is discussed in Sec. VII-C.
@@
Next, by utilizing the $1$-Wasserstein distance and following a convex reformulation of the DR-CBF in [33, Proposition 6.5], lead to the following optimization problem, whose feasibility implies the DRO constraint in (18):
$\mathbf{u}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m},s\in\mathbb{R},\beta_{i}\in\mathbb{R}}\|\mathbf{u}-\bar{\mathbf{u}}(\mathbf{q},t)\|^{2},$ (20)
@@
VII-C Uncertainty Sample Selection
@@ VIII Evaluation
VIII-A Baselines and Parameters
@@
The baseline planners are implemented using the Open Motion Planning Library (OMPL) [40], ensuring consistency across all algorithms. All controllers are implemented using Casadi [4], and the CBF-QP and our DR-CBF-QP problems are solved using the interior point optimizer.
@@
Parameters: Tables I and II summarize the parameters used for the planners and controllers.
TABLE I: Planner parameters used for evaluation. Goal bias specifies the probability of sampling configurations directly toward the goal. Step size defines the maximum extension distance during tree expansion. Safety margin indicates the minimum clearance required for collision checking, and collision check resolution determines the granularity of edge validity checks, expressed as a fraction of the configuration space diagonal. For the 2-DoF planar robot, a resolution of 0.01 corresponds to 0.089 radians, while for the 6-DoF xArm, a resolution of 0.002 corresponds to 0.04 radians.
@@
TABLE II: Control parameters used for evaluation. The PD control gains $K_{P}$ and $K_{D}$ are positive definite matrices, assumed to be diagonal with identical values along the diagonal. The function $\alpha_{h}$ is a linear class $\mathcal{K}_{\infty}$ function used in the CBF formulation. Parameters $k$ and $\zeta$ define the governor dynamics in (14). The risk tolerance $\epsilon$ and sample size $N$ correspond to the DR-CBF in (18), while $r$ represents the Wasserstein radius.
@@
VIII-B 2-DoF Planar Robot Simulation
@@
TABLE III: Planning performance comparison on a 2-DoF planar robot.
Planner Col. Checks Path Length
CDF-RRT $2108.9\pm 438.7$ $3.64\pm 0.69$
SDF-RRT [46] $2118.3\pm 388.9$ $3.61\pm 0.69$
SDF-RRT-Connect $2144.6\pm 459.7$ $3.63\pm 0.74$
SDF-Lazy-RRT $1759.2\pm 299.3$ $3.75\pm 0.83$
Bubble-CDF $\boldsymbol{153.8\pm 62.2}$ $3.72\pm 0.81$
@@
Results and Discussion: Table III summarizes the performance metrics across all 500 environments. The results demonstrate the significant efficiency of the bubble-CDF planner in reducing the number of collision checks while maintaining comparable path quality. The bubble-CDF planner requires $153.8\pm 62.2$ collision checks, while even the best baseline (Lazy-RRT) requires at least 10×10$\times$10 × more collision checks. This showcases the ability of our proposed bubble-CDF planner to explore the configuration space more efficiently. Despite this reduction, the bubble-based planner produces paths with similar lengths ($3.72\pm 0.81$) radians to those of the baselines.
TABLE IV: Control performance comparison on a 2-DoF planar robot.
@@
Results and Discussion: Table IV presents the control performance across static and dynamic environments. The proposed DR-CBF-QP formulation achieves a $100\%$ success rate in static environments and high success rate ($99.2\%$) in dynamic environments, outperforming both the PD controller and the standard CBF-QP approach. The PD controller, designed solely for trajectory tracking without obstacle avoidance, exhibits the lowest success rates in both static and dynamic environments. The baseline CBF-QP shows good performance in static scenarios, but its success rate significantly deteriorates in dynamic environments ($63.8\%$). We attribute this degradation to disregarding uncertainties in dynamic obstacle velocities.
@@
VIII-C 6-DoF xArm Robot Simulation
@@
TABLE V: Planning performance comparison on a 6-DoF xArm robot in Pybullet Simulation.
Planner Col. Checks Path Length
CDF-RRT $3415.4\pm 803.6$ $3.11\pm 0.46$
SDF-RRT [46] $3251.8\pm 782.4$ $3.07\pm 0.49$
SDF-RRT-Connect $3437.4\pm 723.2$ $3.11\pm 0.43$
SDF-Lazy-RRT $2741.1\pm 737.9$ $3.15\pm 0.53$
Bubble-CDF $\boldsymbol{278.5\pm 80.7}$ $3.04\pm 0.46$
@@
Results and Discussion: Table V summarizes the planning results for the 6-DoF xArm robot. Similar to the 2-DoF case, the bubble-CDF planner significantly outperforms baseline methods in terms of collision checks. Path lengths across all planners remain comparable, indicating that the bubble-CDF approach does not compromise solution quality.
@@
Table VI quantitatively compares the control performance of different controllers under both static and dynamic environments. The baseline PD controller achieves a success rate of 68% in the static environment but struggles significantly in dynamic scenarios, with a success rate dropping to 18%, highlighting its inability to effectively react to moving obstacles.
@@
TABLE VI: Control performance comparison on a 6-DoF xArm robot in Pybullet Simulation.
@@
VIII-D 6-DoF xArm Robot Experiments
TABLE VII: Planning performance comparison on a real 6-DoF xArm robot.
Planner Col. Checks Path Length
CDF-RRT $5617.4\pm 1333.6$ $3.71\pm 0.71$
SDF-RRT [46] $5911.9\pm 1487.7$ $3.72\pm 0.68$
SDF-RRT-Connect $5801.5\pm 1265.4$ $3.75\pm 0.63$
SDF-Lazy-RRT $4765.8\pm 1331.4$ $3.84\pm 0.69$
Bubble-CDF $\boldsymbol{634.8\pm 225.1}$ $3.81\pm 0.66$
@@
Results and Discussion: Table VII presents a quantitative comparison of planning performance across different methods. The bubble-CDF planner significantly reduces the number of collision checks compared to all baselines, achieving a nearly tenfold reduction. This improvement stems from the use of configuration-space bubbles, which certify local collision-free regions and minimize the need for frequent edge-based collision checking. Consistent with simulation results, the generated path lengths remain comparable across all methods, demonstrating that our approach maintains solution quality while substantially improving efficiency.
@@
IX Limitations and Future Directions
While our proposed bubble-CDF planner and DR-CBF safety filter demonstrate strong performance in motion planning and control for robotic manipulators, several limitations demand further investigation. First, a fundamental challenge lies in balancing safety and stability, particularly in highly cluttered environments. While the DR-CBF-QP formulation ensures collision-free execution, it does not explicitly account for cases where the robot may become stuck in front of obstacles, unable to progress toward the planned path. This incompatibility between obstacle avoidance and goal-reaching poses a significant challenge, and future work should focus on addressing the compatibility of safety and stability.
Second, although our formulation is designed to handle uncertainty in CDF modeling, point-cloud observations, and dynamic obstacle velocity estimates, reliable obstacle point clouds and velocity estimation remains important for real-world deployments. In real-world experiments, we currently rely on ArUco markers for velocity tracking, but a more general and robust approach, such as learning-based perception or multi-frame point-cloud tracking, could improve adaptability in unstructured environments.
Third, the current framework does not explicitly consider self-collision, which may lead to infeasible or unsafe configurations where the robot intersects with itself. Incorporating self-collision constraints into the neural CDF representation or integrating explicit collision-checking mechanisms within the motion planner would enhance the framework’s applicability to more complex robotic systems.
Finally, the use of configuration-space bubbles for planning, while computationally efficient, may be overly conservative in densely cluttered environments. Since spherical safe regions do not fully exploit free space, alternative representations such as ellipsoidal or polyhedral safe regions could enable more flexible and efficient motion planning.
X Conclusion

### v2

@@ VI-A Governor-based Trajectory Tracking
VI-B Distributionally Robust Neural CDF Barrier
@@
where $\alpha_{h}$ is a class ${\cal K}_{\infty}$ function.
@@
$\mathbf{u}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{u}\in\mathcal{U}}\|\mathbf{u}-\bar{\mathbf{u}}(\mathbf{q},t)\|^{2}$ (21)
s.t. $\;\text{CBC}(\mathbf{q},\mathbf{u},t)\geq 0,$
where ${\cal U}:=\{\mathbf{u}_{\min}\leq\mathbf{u}\leq\mathbf{u}_{\max}\}$ defines the admissible control space, ensuring the commanded manipulator velocities remain within feasible operational limits. However, the CDF barrier $h(\mathbf{q},t)$ and its gradients are not known exactly but rather are approximated through the neural CDF barrier $\hat{h}(\mathbf{q},t)$ in (12). Due to uncertainties introduced by the neural network model and the point-cloud data, $\hat{h}$ cannot be treated as an exact representation of $h$. Instead, we work with samples of the associated uncertainty vector $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$, defined in (20), which capture the variations in $\hat{h}$, its gradient, and its time derivative. The practical approach of obtaining such samples is discussed in Sec. VI-C.
@@
Next, by utilizing the $1$-Wasserstein distance and following a convex reformulation of the DR-CBF in [34, Proposition 6.5], lead to the following optimization problem, whose feasibility implies the DRO constraint in (22):
$\mathbf{u}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{u}\in{\cal U},s\in\mathbb{R},\beta_{i}\in\mathbb{R}}\|\mathbf{u}-\bar{\mathbf{u}}(\mathbf{q},t)\|^{2},$ (24)
@@
VI-C Uncertainty Sample Selection
@@ VII Evaluation
VII-A Implementation Details
Training Dataset and Loss Function: We discretize the robot’s workspace into a $40\times 40\times 40$ grid. For each workspace point, we generate contact configurations using BFGS optimization, and then select 200 diverse configurations via farthest-point sampling in configuration space [29]. During training, we employ an online batching strategy where each batch randomly samples 10 configurations and 50 workspace points. For each configuration–point pair $(\mathbf{p}_{i},\mathbf{q}_{i})$, the ground-truth environment CDF value $f_{c}^{\star}(\mathbf{p}_{i},\mathbf{q}_{i})$ is computed as the minimum joint-space distance to the set of contact configurations associated with $\mathbf{p}_{i}$.
The network is trained using the following loss function:
$\mathcal{L}_{\text{env}}=\frac{1}{N_{1}N_{2}}\sum_{i=1}^{N_{1}}\sum_{j=1}^{N_{2}}\Big{(}\hat{f}_{c}(\mathbf{p}_{i},\mathbf{q}_{j};\boldsymbol{\theta}_{1})-f_{c}^{\star}(\mathbf{p}_{i},\mathbf{q}_{j})\Big{)}^{2}\quad$
$+\lambda\left(\|\nabla_{\mathbf{q}}\hat{f}_{c}(\mathbf{p}_{i},\mathbf{q}_{j};\boldsymbol{\theta}_{1})\|-1\right)^{2},$ (27)
where $\lambda=0.05$ is a user-specified weight.
For the self-collision CDF (SCDF), we generate training data by uniformly sampling joint configurations and detecting self-collisions using the PyBullet simulation engine [8]. Each self-colliding configuration $\mathbf{q}^{\prime}\in\mathcal{C}_{\mathrm{sc}}$ is stored along with the minimal joint index range $[a:b]$ responsible for the collision. During training, each batch consists of $N$ randomly sampled configurations $\{\mathbf{q}_{j}\}_{j=1}^{N}$. For each configuration, the ground-truth SCDF value $f_{\mathrm{sc}}^{\star}(\mathbf{q}_{j})$ is computed according (8).
The SCDF network is trained with the following loss:
$\mathcal{L}_{\text{sc}}=\frac{1}{N_{3}}\sum_{j=1}^{N_{3}}\Big{(}\hat{f}_{\mathrm{sc}}(\mathbf{q}_{j};\boldsymbol{\theta}_{2})-f_{\mathrm{sc}}^{\star}(\mathbf{q}_{j})\Big{)}^{2}\quad$
$+\lambda\left(\|\nabla_{\mathbf{q}}\hat{f}_{\mathrm{sc}}(\mathbf{q}_{j};\boldsymbol{\theta}_{2})\|-1\right)^{2},$ (28)
where $\lambda=0.05$ is user-specified weight.
Network Architecture and Training Setup: The neural CDF and SCDF are each implemented as a 5-layer fully connected multilayer perceptron (MLP) with GELU activation functions [21]. Skip connections are included at the second and fourth layers to improve learning stability. Both networks are trained using PyTorch on an NVIDIA RTX 4090 GPU for 50,000 epochs with the Adam optimizer and a learning rate of $2\times 10^{-4}$.
Training Results: The environment CDF model is evaluated on a held-out test set consisting of 50,000 triplets of configurations ($\mathbf{q}$), workspace points ($\mathbf{p}$), and corresponding CDF values. It achieves a Mean Absolute Error (MAE) of 0.041 radians and a Root Mean Square Error (RMSE) of 0.073 radians. Similarly, the SCDF model is evaluated on 10,000 held-out configurations with known self-collision distances, achieving MAE of 0.029 radians and RMSE of 0.054 radians. As shown in Table I, the combined inference time for evaluating both the environment CDF and the self-collision CDF, which together define the neural CDF barrier in (12), remains efficient across different batch sizes. This scalability allows the framework to be effectively integrated with sampling-based motion planning algorithms and real-time control using point cloud observations.
TABLE I: Inference time (ms) of Neural CDF barrier on xArm for varying batch sizes of configurations ($\mathbf{q}$) and workspace points ($\mathbf{p}$).
$\mathbf{q}$ $\mathbf{p}$ 1 10 100 1000 10000
1 0.272 0.281 0.295 0.291 0.742
10 0.281 0.307 0.297 0.758 6.818
100 0.301 0.301 0.763 7.060 70.215
VII-B Baselines and Parameters
@@
The baseline planners are implemented using the Open Motion Planning Library (OMPL) [41], ensuring consistency across all algorithms. All controllers are implemented using Casadi [4], and the CBF-QP and our DR-CBF-QP problems are solved using the interior point optimizer, operating at a control frequency of $50$ Hz.
@@
Parameters: Tables II and III summarize the parameters used for the planners and controllers.
TABLE II: Planner parameters used for evaluation. Goal bias specifies the probability of sampling configurations directly toward the goal. Step size defines the maximum extension distance during tree expansion. Safety margin indicates the minimum clearance required for collision checking, and collision check resolution determines the granularity of edge validity checks, expressed as a fraction of the configuration space diagonal. For the 2-DoF planar robot, a resolution of 0.01 corresponds to 0.089 radians, while for the 6-DoF xArm, a resolution of 0.002 corresponds to 0.04 radians.
@@
TABLE III: Control parameters used for evaluation. The PD control gains $K_{P}$ and $K_{D}$ are positive definite matrices, assumed to be diagonal with identical values along the diagonal. The function $\alpha_{h}$ is a linear class $\mathcal{K}_{\infty}$ function used in the CBF formulation. Parameters $k$ and $\zeta$ define the governor dynamics in (18). The risk tolerance $\epsilon$ and sample size $N$ correspond to the DR-CBF in (22), while $r$ represents the Wasserstein radius.
@@
VII-C 2-DoF Planar Robot Simulation
@@
TABLE IV: Planning performance comparison on a 2-DoF planar robot.
Planner Col. Checks Path Length Time (s)
CDF-RRT $2108.9\pm 438.7$ $3.64\pm 0.69$ $0.33\pm 0.06$
SDF-RRT [47] $2118.3\pm 388.9$ $3.61\pm 0.69$ $0.29\pm 0.04$
SDF-RRT-Connect $2144.6\pm 459.7$ $3.63\pm 0.74$ $0.33\pm 0.05$
SDF-Lazy-RRT $1759.2\pm 299.3$ $3.75\pm 0.83$ $0.31\pm 0.06$
Bubble-CDF $\boldsymbol{153.8\pm 62.2}$ $3.72\pm 0.81$ $\boldsymbol{0.13\pm 0.04}$
@@
Results and Discussion: Table IV summarizes the performance metrics across all 500 environments. All planners succeed in finding the path. The results demonstrate the significant efficiency of the bubble-CDF planner in reducing the number of collision checks, reducing the planning time, while maintaining comparable path quality. The bubble-CDF planner requires $153.8\pm 62.2$ collision checks, while even the best baseline (Lazy-RRT) requires at least 10×10$\times$10 × more collision checks. As a result, bubble-CDF achieves the lowest planning time of all methods at $0.17\pm 0.05$ seconds. This showcases the ability of our proposed bubble-CDF planner to explore the configuration space more efficiently. Despite this reduction, the bubble-based planner produces paths with similar lengths ($3.72\pm 0.81$) radians to those of the baselines.
TABLE V: Control performance comparison on a 2-DoF planar robot.
@@
Results and Discussion: Table V presents the control performance across static and dynamic environments. The proposed DR-CBF-QP formulation achieves a $100\%$ success rate in static environments and high success rate ($99.2\%$) in dynamic environments, outperforming both the PD controller and the standard CBF-QP approach. The PD controller, designed solely for trajectory tracking without obstacle avoidance, exhibits the lowest success rates in both static and dynamic environments. The baseline CBF-QP shows good performance in static scenarios, but its success rate significantly deteriorates in dynamic environments ($63.8\%$). We attribute this degradation to disregarding uncertainties in dynamic obstacle velocities.
@@
VII-D 6-DoF xArm Robot Simulation
@@
TABLE VI: Planning performance comparison on a 6-DoF xArm robot in Pybullet Simulation.
Planner Col. Checks Path Length Time (s)
CDF-RRT $3415.4\pm 803.6$ $3.11\pm 0.46$ $0.75\pm 0.23$
SDF-RRT [47] $3251.8\pm 782.4$ $3.07\pm 0.49$ $0.73\pm 0.20$
SDF-RRT-Connect $3437.4\pm 723.2$ $3.11\pm 0.43$ $0.71\pm 0.21$
SDF-Lazy-RRT $2741.1\pm 737.9$ $3.15\pm 0.53$ $0.72\pm 0.22$
Bubble-CDF $\boldsymbol{278.5\pm 80.7}$ $3.04\pm 0.46$ $\boldsymbol{0.15\pm 0.06}$
@@
Results and Discussion: Table VI summarizes the planning results for the 6-DoF xArm robot. Similar to the 2-DoF case, the bubble-CDF planner significantly outperforms baseline methods in terms of collision checks and planning time. Path lengths across all planners remain comparable, indicating that the bubble-CDF approach does not compromise solution quality.
@@
Table VII quantitatively compares the control performance of different controllers under both static and dynamic environments. The baseline PD controller achieves a success rate of 68% in the static environment but struggles significantly in dynamic scenarios, with a success rate dropping to 18%, highlighting its inability to effectively react to moving obstacles.
@@
TABLE VII: Control performance comparison on a 6-DoF xArm robot in Pybullet Simulation.
@@
VII-E 6-DoF xArm Robot Experiments
TABLE VIII: Planning performance comparison on a real 6-DoF xArm robot.
Planner Col. Checks Path Length Time (s)
CDF-RRT $5617.4\pm 1333.6$ $3.71\pm 0.71$ $1.02\pm 0.33$
SDF-RRT [47] $5911.9\pm 1487.7$ $3.72\pm 0.68$ $1.04\pm 0.33$
SDF-RRT-Connect $5801.5\pm 1265.4$ $3.75\pm 0.63$ $1.05\pm 0.36$
SDF-Lazy-RRT $4765.8\pm 1331.4$ $3.84\pm 0.69$ $0.96\pm 0.32$
Bubble-CDF $\boldsymbol{634.8\pm 225.1}$ $3.81\pm 0.66$ $\boldsymbol{0.22\pm 0.08}$
@@
Results and Discussion: Table VIII presents a quantitative comparison of planning performance across different methods. The proposed bubble-CDF planner achieves an order-of-magnitude reduction in the number of collision checks compared to all baseline methods and reduces the average planning time by a factor of five, decreasing from 1.0 seconds to about 0.2 seconds. This improvement stems from the use of configuration-space bubbles, which certify local collision-free regions and minimize the need for frequent edge-based collision checking. Consistent with simulation results, the generated path lengths remain comparable across all methods, demonstrating that our approach maintains solution quality while substantially improving efficiency.
@@
VIII Conclusion